OpenStudy (anonymous):

Find the slope of a line perpendicular to 2x-y=16 8 2 1/2 -1/2 * what's the formula ? So I can try and figure it out first

OpenStudy (nurali):

well first you must find the slope originally: 2x=16+y 2x-16=y therefore the slope is 2 so perpendicular means put this value under 1 and take the negative of it: so your slope is 1/2 and then take the negative: -1/2 is your answer

OpenStudy (triciaal):

what's the formula ? So I can try and figure it out first if lines are perpendicular the product of the slopes is -1 m* m1 = -1 slope of line 1 = m slope of perpendicular, m1 = -1/m